# Uncertainty-Aware PPG-2-ECG for Enhanced Cardiovascular Diagnosis using Diffusion Models

### v1

@@ 1 Introduction
Cardiovascular diseases are a significant public health problem, affecting millions of people worldwide and remaining a leading cause of mortality [1]. The analysis of cardiovascular conditions using Electrocardiography (ECG) signals has emerged as one of the most widely utilized diagnostic tool [2], and its practice and efficiency have been continuously refined over many decades.
@@
While ECG sensing is non-invasive and relatively easy to acquire, it is time-consuming and requires the expertise of trained professionals with specialized skills in order to ensure accurate diagnosis. Consequently, there has been a surge in the development of numerous contemporary wearable ECG systems in recent decades for digital diagnosis. However, the materials used to deliver a high-quality signal via electrodes can often lead to skin irritation and discomfort during extended usage, thereby limiting the long-term viability of these devices [3].
@@
A possible alternative to ECG sensing and diagnosis is the use of Photoplethysmography (PPG), which is an optically-based non-invasive signal associated with rhythmic changes in blood volume within tissues [4]. Unlike ECG, PPG signals are easier to obtain, convenient, and cost-effective, as they are widely available in clinics and hospitals and can be sensed through finger/toe clips. Moreover, the popularity of PPG is increasing with the emergence of wearable devices like smart watches, enabling continuous long-term monitoring without causing skin irritations.
@@
Naturally, the PPG and ECG signals are inter-related, as the timing, amplitude, and shape characteristics of the PPG waveform contain information about the interaction between the heart and the blood vessels. These features of the PPG have been leveraged to measure heart rate, heart rate variability, respiration rate [5], blood oxygen saturation [6], blood pressure [7], and to assess vascular function [8, 9]. Despite these capabilities, using PPG-based monitoring during daily activities and light physical exercises poses unavoidable uncertainty challenges due to signal inaccuracies and inherent noise, stemming from their indirect nature, compared to ECG signals sourced directly from the heart.
@@
As the use of wearable devices capturing PPG signals is becoming widespread, there is an emerging interest in accurately converting PPG sources into ECG signals [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, such a mapping can be regarded as an ill-posed inverse problem. This implies that multiple ECG solutions may correspond to the same input PPG signal. Nevertheless, previous work in this domain has tended to focus on providing a single ECG solution to the conversion task, ignoring the inherent uncertainty in the conversion and thus failing to capture the complete information contained in the PPG signal. Such studies suggest optimizing neural models either via regularized regression with an $L_{1}$/$L_{2}$ loss [11, 10, 12, 15, 16, 17, 19, 20, 21] or relying on a generative approach that produces a single solution that approximates a sample from the posterior distribution of ECG given the PPG signal [13, 14, 18, 22]. Here, we argue that both these strategies are lacking, and the conversion in question calls for a different approach.
@@
In this paper, we introduce Uncertainty-Aware PPG-2-ECG (UA-P2E), a novel PPG-2-ECG conversion methodology, aimed at enhancing the classification of cardiovascular conditions using ECG signals derived from PPG data. Unlike prior work, our approach aims to address both the uncertainty arising from the conversion task, regarding the spread and variability of the possible ECG solutions given the PPG [23, 24], and also the uncertainty in the classification task, enabling the selection of PPG samples with low classification risk and high confidence [25].
@@
To achieve the above, we leverage a conditional diffusion-based methodology [26, 27, 28]. These generative models have recently emerged as the leading synthesis approach in various tasks [29, 30], and have been shown to enable high-fidelity posterior modeling for inverse problems [30, 31, 32, 33, 34]. We utilize this capability to produce a multitude of posterior sampled ECG results that correspond to the input PPG, and provide classification decisions based on this cluster of outcomes. Additionally, we introduce several methods for displaying the resulting ECG candidates for better human interpretability. Figure 1 provides an illustration of our approach to the conversion-classification process.
@@
Our work presents a thorough empirical study demonstrating superior performance in classification and uncertainty estimation compared to baseline methods. Our main results are demonstrated on the “Computing in Cardiology” (CinC) dataset [35], which provides the most challenging classification data in cardiology.
@@ 2 Related Work
Naturally, our work is not the first one to consider the conversion of PPG signals to ECG ones. Below we outline the main trends in this line of work, and add a discussion on uncertainty quantification, as we aim to practice in this paper.
Regression: Earlier approaches to the PPG-2-ECG conversion problem primarily focused on approximating either the Minimum Mean Absolute Error (MMAE) estimator using $L_{1}$ loss [11, 10, 16, 19] or the Minimum Mean Squared Error (MMSE) estimator with $L_{2}$ loss [12, 15, 17, 20, 21]. These works commonly employed regularization techniques to enhance perceptual outcomes. However, the work reported in [36] which discovered the distortion-perception tradeoff has shown that when optimizing for distortion (of any kind), perceptual quality is necessarily compromised, thus pushing the solutions in the above methods to signals that are likely to drift out of the valid ECG manifold, thus weakening their classification performance.
Generation: The first attempt at PPG-2-ECG conversion using generative models was introduced by the authors of CardioGAN [13]. They employed an adversarial training scheme with an encoder-decoder architecture. Similarly, P2E-WGAN [14] utilized conditional Wasserstein GANs for this task. [18] introduced a variational inference methodology to approximate the posterior distribution with a Gaussian prior assumption. The authors of RDDM [22] have first adapted a diffusion-based model that accelerates a DDPM [27] diffusion process for this conversion task. However, despite the success of the generative approach, all the above works have chosen to use a single (possibly random) ECG solution from the posterior distribution, thus missing the complete picture of the possibilities of the different ECGs given the PPG signal.
Uncertainty Quantification: In the framework of the conversion-classification explored in this paper, uncertainty quantification serves a dual purpose. Firstly, it seeks to characterize the range, spread, and variability of potential solutions for a given input PPG signal [23, 24]. Secondly, it also aims at assessing the confidence in the correctness of classification predictions, e.g. via a selective classification [25] scheme, which we adopt in this paper. Building upon principles from these two threads of work, we characterize the diverse solutions that can potentially describe the input, leading to an improved methodology within the conversion-classification framework.
@@ 3 Problem Formulation and Goals
Following the data processing inequality [37], any transformation $\hat{X}=\mathbb{E}_{Z^{\prime}\sim\pi(Z^{\prime})}[g(Y,Z^{\prime})]$ satisfies the inequality
@@
Intuitively, this strategy averages the classification scores of posterior samples that emerge from the posterior distribution $\pi(X|Y)$, rather than averaging the samples themselves. In Appendix A we prove Theorem A.1, implying that our approach utilizes the full extent of the information within the observation $Y$, and is therefore optimal.
@@ 4 UA-P2E: Uncertainty-Aware PPG-2-ECG
While our proposed approach is applicable using any conditional stochastic generative estimator, denoted as $g:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathcal{X}$, we focus in this work on conditional diffusion-based models [26, 27, 28], which have emerged as the leading synthesis approach in various domains [29, 30]. Our diffusion process is trained and evaluated on raw signal data of fixed-length $Y,X\in\mathbb{R}^{d}$ PPG and ECG signal pairs, sampled at $125$Hz, thus covering $d/125$ seconds.111in our experiments $d=1024$ and thus the signals cover $\approx 8$ seconds. The central part of the diffusion process involves a conditional denoiser, denoted as $\epsilon(X_{t},Y,t)$. This denoiser is trained to predict the noise in $X_{t}$, which is a combination of $X_{0}$ (representing a clean instance) and white additive Gaussian noise at a noise level denoted by $t$. In this conditional process, the denoiser leverages the PPG signal $Y$ and the noise level $t$ as supplementary information. Training $\epsilon$ is based on pairs of $(X,Y)$ signals - a clean ECG and its corresponding PPG. After training, we apply multiple diffusion operations on a given PPG signal $Y\in\mathcal{Y}$ to draw potential ECG solutions $\{\hat{X}_{i}\}_{i=1}^{K}$, where each $\hat{X}_{i}\sim\hat{\pi}(X|Y)$ is a sample from the approximated posterior distribution of the ECG given the PPG signal.
In Section 4.1 we describe our enhanced classification procedure that relies on the above posterior sampler, denoted by $g$. Additionally, in Section 4.2 we discuss how to improve the quantification of uncertainty in classification, aiming to enhance confidence in predictions. Lastly, in Section 4.3, we offer methods to visualize the cloud of ECG solutions $\{\hat{X}_{i}\}_{i=1}^{K}$. This visualization aims at enhancing interpretability for medical professionals, allowing for a better understanding of the various ECG signals that emerge from a given PPG signal.
@@ 4.1 UA-P2E: Enhanced Classification
The methodology for UA-P2E’s enhanced classification consists of three phases. The first involves training the probabilistic classification model, denoted as $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{+}$, using ECG signals and their associated labels. $f_{\mathcal{X}}$ can be optimized through the commonly used Cross-Entropy loss function, producing an output that has a probabilistic interpretation. In the second phase, each potential posterior sample $\hat{X}_{i}$ from the set $\{\hat{X}_{i}\}_{i=1}^{K}$ (derived by applying $g$ to a given $Y$) is evaluated by this classifier, extracting an output score $f_{\mathcal{X}}(\hat{X}_{i})$, being an estimated probability of belonging to the positive class ($C=1$). In the final phase, the average classification score, $\frac{1}{K}\sum_{i=1}^{K}f_{\mathcal{X}}(\hat{X}_{i})$, is computed, and compared to a decision threshold $t\in\mathbb{R}^{+}$. The complete workflow of our proposed method to enhance classification is detailed in Algorithm 1.
@@
Algorithm 1 seeks to approximate the ESC (Definition 3.2), which has been proven optimal in Theorem A.1.
4.2 UA-P2E: Uncertainty Quantification
The proposed PPG-2-ECG conversion has two layers of uncertainty awareness, as described herein. The first takes into account the inherent uncertainty that emerges from the ill-posed conversion itself, quantifying the spread and variability of the possible ECG candidates that derive from the PPG inputs. In Appendix C, we present empirical evidence illustrating the uncertainty from this perspective. The second uncertainty layer leverages the above property for providing classification predictions with a better confidence in their correctness. This serves a selective classification [25] scheme, in which a carefully chosen subset of the test PPG data is selected by their calculated confidence, exhibiting improved accuracy.
Selective classification [25] enables handling cases in which classifying every PPG signal within a test set is challenging due to uncertainty or data corruption. This technique allows for abstaining from making predictions for signals with low confidence, thus reducing classification errors on the remaining data. A confidence function, defined as $\kappa:\mathcal{Y}\rightarrow\mathbb{R}^{+}$, serves as the engine for this task by predicting the correctness of predictions. One common choice of this confidence function $\kappa$ is to use the maximal softmax score of a classifier. In a similar spirit, in this work we set $\kappa(Y)$ to take the value $\kappa(Y)=\max\{1-f_{\mathcal{Y}}(Y),f_{\mathcal{Y}}(Y)\}$, where $f_{\mathcal{Y}}(Y)$ approximates the ESC (Definition 3.2), as summarized in Algorithm 1.
where $\delta\in(0,1)\ll 1$ is a user-defined calibration error. Intuitively, by setting $\alpha=\delta=0.1$, the above implies that the classification error among selected PPG signals, whose confidence scores are greater than $\lambda$, will be lower than $\alpha$ ($=10\%$) with probability of $1-\delta$ ($=90\%$). In Appendix B, we describe how to calibrate $\lambda$ to provide the guarantee stated in Equation (2).
@@ 4.3 Visualizing UA-P2E’s ECG Output(s)
In Appendix D, we provide empirical results regarding the quality of the three visualization approaches mentioned above. These results suggest that the most likely score ECG, as defined in Equation (3), provides the highest quality.
@@ 5 Empirical Study
Unlike prior work, this paper aims to address the uncertainty in the conversion of PPG to ECG signals, and this is obtained by leveraging a diffusion-based conversion methodology. More specifically, we adopt the DDIM framework [28], aiming to reduce the number of iterations (in our experiments we applied $T=100$ iterations) required for the sampling procedure, and thus increasing the efficiency of our conversion algorithm. The diffusion denoising model is trained using the MIMIC-III matched waveform database [38], which is one of the most extensive dataset for paired PPG and ECG signals.
@@
As for the classification objective, we train and evaluate the classification model, denoted as $f_{\mathcal{X}}$, using the ‘Computing in Cardiology” (CinC) [35] dataset, containing multi-labeled ECG signals of 11 cardiovascular conditions  [^2] The following cardiovascular conditions are considered: (1) Ventricular premature beats, (2) Atrial flutter, (3) Sinus rhythm, (4) Atrial fibrillation, (5) Supraventricular premature beats, (6) Bradycardia, (7) Pacing rhythm, (8) Sinus tachycardia, (9) 1st degree AV block, (10) Sinus bradycardia, and (11) Sinus arrhythmia. that are detectable by lead II ECG signals. Interestingly, our experiments rely in part on a reversed diffusion-based conversion model, designed to sample from the likelihood distribution $\hat{\pi}(Y|X)$, i.e. converting ECG (lead II) to PPG. This reversed sampling is utilized in order to complete the CinC [35] dataset used in our tests, enabling an analysis of challenging cardiac classification tasks. Figure 2 (bottom row) presents typical sampling outcomes of the reversed conversion model.
For further details on experimental details, including the datasets used for training and evaluating the conversion and classification models, along with details on data pre-processing and training schemes, we refer the reader to Appendix E.

### v3

@@ 1 Introduction
Cardiovascular diseases are a significant public health problem, affecting millions of people worldwide and remaining a leading cause of mortality (Tsao et al., 2022). The analysis of cardiovascular conditions using Electrocardiography (ECG) signals has emerged as one of the most widely utilized diagnostic tool (Kligfield et al., 2007), and its practice and efficiency have been continuously refined over many decades.
@@
While ECG sensing is non-invasive and relatively easy to acquire, it is time-consuming and requires the expertise of trained professionals with specialized skills in order to ensure accurate diagnosis. Consequently, there has been a surge in the development of numerous contemporary wearable ECG systems in recent decades for digital diagnosis. However, the materials used to deliver a high-quality signal via electrodes can often lead to skin irritation and discomfort during extended usage, thereby limiting the long-term viability of these devices (Zhu et al., 2019).
@@
A possible alternative to ECG sensing and diagnosis is the use of Photoplethysmography (PPG), which is an optically-based non-invasive signal associated with rhythmic changes in blood volume within tissues (Reisner et al., 2008). Unlike ECG, PPG signals are easier to obtain, convenient, and cost-effective, as they are widely available in clinics and hospitals and can be sensed through finger/toe clips. Moreover, the popularity of PPG is increasing with the emergence of wearable devices like smart watches, enabling continuous long-term monitoring without causing skin irritations.
@@
Naturally, the PPG and ECG signals are inter-related, as the timing, amplitude, and shape characteristics of the PPG waveform contain information about the interaction between the heart and the blood vessels. These features of the PPG have been leveraged to measure heart rate, heart rate variability, respiration rate (Karlen et al., 2013), blood oxygen saturation (Aoyagi & Miyasaka, 2002), blood pressure (Payne et al., 2006), and to assess vascular function (Marston, 2002; Allen & Murray, 1993). Despite these capabilities, using PPG-based monitoring during daily activities and light physical exercises poses unavoidable uncertainty due to signal inaccuracies and inherent noise, stemming from their indirect nature, compared to ECG signals sourced directly from the heart.
@@
As the use of wearable devices capturing PPG signals is becoming widespread, there is an emerging interest in accurately converting PPG sources into ECG signals (Chiu et al., 2020; Li et al., 2020; Zhu et al., 2021; Sarkar & Etemad, 2021; Vo et al., 2021; Tang et al., 2022; Ho et al., 2022; Tian et al., 2022; Vo et al., 2023; Lan, 2023; Abdelgaber et al., 2023; Hinatsu et al., 2023; Shome et al., 2023). However, such a mapping can be regarded as an ill-posed inverse problem. This implies that multiple ECG solutions may correspond to the same input PPG signal. Nevertheless, previous work in this domain has tended to focus on providing a single ECG solution to the conversion task, ignoring the inherent uncertainty in the conversion and thus failing to capture the complete information contained in the PPG signal. Such studies suggest optimizing neural models either via regularized regression with an $L_{1}$/$L_{2}$ loss (Li et al., 2020; Chiu et al., 2020; Zhu et al., 2021; Tang et al., 2022; Ho et al., 2022; Tian et al., 2022; Lan, 2023; Abdelgaber et al., 2023; Hinatsu et al., 2023) or relying on a generative approach that produces a single solution that approximates a sample from the posterior distribution of ECG given the PPG signal (Sarkar & Etemad, 2021; Vo et al., 2021; 2023; Shome et al., 2023). Here, we argue that both these strategies are lacking, and the conversion in question calls for a different approach.
@@
In this paper, we introduce Uncertainty-Aware PPG-2-ECG (UA-P2E), a novel PPG-2-ECG conversion methodology, aimed at enhancing the classification of cardiovascular conditions using ECG signals derived from PPG data. Unlike prior work, our approach aims to address both the uncertainty arising from the conversion task, regarding the spread and variability of the possible ECG solutions given the PPG (Angelopoulos et al., 2022; Belhasin et al., 2023), and also the uncertainty in the classification task, enabling the selection of PPG samples with low classification risk and high confidence (Geifman & El-Yaniv, 2017).
@@
To achieve the above, we leverage a conditional diffusion-based methodology (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020). These generative models have recently emerged as the leading synthesis approach in various tasks (Dhariwal & Nichol, 2021; Yang et al., 2023), and have been shown to enable high-fidelity posterior modeling for inverse problems (Yang et al., 2023; Kawar et al., 2021; 2022; Song et al., 2022; Chung et al., 2022). We utilize this capability to produce a multitude of posterior sampled ECG results that correspond to the input PPG, and provide classification decisions based on this cluster of outcomes. Additionally, we introduce several methods for displaying the resulting ECG candidates for better human interpretability. Figure 1 provides an illustration of our approach to the conversion-classification process.
@@
Our work presents a thorough empirical study demonstrating superior performance in classification and uncertainty estimation compared to baseline methods. Our main results are demonstrated on the “Computing in Cardiology” (CinC) dataset (Reyna et al., 2021), which provides the most challenging classification data in cardiology.
@@ 2 Related Work
Naturally, our work is not the first one to consider the conversion of PPG signals to ECG ones. Below we outline the main trends in this line of work.
Regression: Earlier approaches to the PPG-2-ECG conversion problem primarily focused on approximating either the Minimum Mean Absolute Error (MMAE) estimator using $L_{1}$ loss (Li et al., 2020; Chiu et al., 2020; Ho et al., 2022; Lan, 2023) or the Minimum Mean Squared Error (MMSE) estimator with $L_{2}$ loss (Zhu et al., 2021; Tang et al., 2022; Tian et al., 2022; Abdelgaber et al., 2023; Hinatsu et al., 2023). These works commonly employed regularization techniques to enhance perceptual outcomes. However, the work reported in Blau & Michaeli (2018) which discovered the distortion-perception tradeoff has shown that when optimizing for distortion (of any kind), perceptual quality is necessarily compromised, thus pushing the solutions in the above methods to signals that are likely to drift out of the valid ECG manifold, weakening their classification performance.
Generation: The first attempt at PPG-2-ECG conversion using generative models was introduced by the authors of CardioGAN (Sarkar & Etemad, 2021). They employed an adversarial training scheme with an encoder-decoder architecture. Similarly, P2E-WGAN (Vo et al., 2021) utilized conditional Wasserstein GANs for this task. Vo et al. (2023) introduced a variational inference methodology to approximate the posterior distribution with a Gaussian prior assumption. The authors of RDDM (Shome et al., 2023) have first adapted a diffusion-based model that accelerates a DDPM (Ho et al., 2020) diffusion process for this conversion task. However, despite the success of the generative approach, all the above works have chosen to use a single (possibly random) ECG solution from the posterior distribution, thus missing the complete picture of the possibilities of the different ECGs given the PPG signal.
@@ 3 Problem Formulation and Goals
Following the data processing inequality (Beaudry & Renner, 2011), any transformation $\hat{X}=\mathbb{E}_{Z^{\prime}\sim\pi(Z^{\prime})}[g(Y,Z^{\prime})]$ satisfies the inequality
@@
Intuitively, this strategy averages the classification scores of posterior samples that emerge from the posterior distribution $\pi(X|Y)$, rather than averaging the samples themselves. As the following Theorem claims, this classification approach is optimal.
Theorem 3.1 (Optimality of the Expected Score Classifier).
  1. 1.
  2. 2.
In Appendix A we prove Theorem 3.1, implying that our approach utilizes the full extent of the information within the observation $Y$, and is therefore optimal.
@@ 4 UA-P2E: Uncertainty-Aware PPG-2-ECG
While our proposed approach is applicable using any conditional stochastic generative estimator, denoted as $g:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathcal{X}$, we focus in this work on conditional diffusion-based models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), which have emerged as the leading synthesis approach in various domains (Dhariwal & Nichol, 2021; Yang et al., 2023). Our diffusion process is trained and evaluated on raw signal data of fixed-length $Y,X\in\mathbb{R}^{d}$ PPG and ECG signal pairs, sampled at $125$Hz, thus covering $d/125$ seconds.111in our experiments $d=1024$ and thus the signals cover $\approx 8$ seconds. The central part of the diffusion process involves a conditional denoiser, denoted as $\epsilon(X_{t},Y,t)$. This denoiser is trained to predict the noise in $X_{t}$, which is a combination of $X_{0}$ (representing a clean instance) and white additive Gaussian noise at a noise level denoted by $t$. In this conditional process, the denoiser leverages the PPG signal $Y$ and the noise level $t$ as supplementary information. Training $\epsilon$ is based on pairs of $(X,Y)$ signals - a clean ECG and its corresponding PPG. After training, we apply multiple diffusion operations on a given PPG signal $Y\in\mathcal{Y}$ to draw potential ECG solutions $\{\hat{X}_{i}\}_{i=1}^{K}$, where each $\hat{X}_{i}\sim\hat{\pi}(X|Y)$ is a sample from the approximated posterior distribution of the ECG given the PPG signal.
@@ 4.1 UA-P2E: Enhanced Classification
The methodology for UA-P2E’s enhanced classification consists of three phases. The first involves training the probabilistic classification model, denoted as $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{+}$, using ECG signals and their associated labels. $f_{\mathcal{X}}$ can be optimized through the commonly used Cross-Entropy loss function, producing an output that has a probabilistic interpretation. In the second phase, each potential posterior sample $\hat{X}_{i}$ from the set $\{\hat{X}_{i}\}_{i=1}^{K}$ (derived by applying $g$ to a given $Y$) is evaluated by this classifier, extracting an output score $f_{\mathcal{X}}(\hat{X}_{i})$, being an estimated probability of belonging to the positive class ($C=1$). In the final phase, the average classification score, $\frac{1}{K}\sum_{i=1}^{K}f_{\mathcal{X}}(\hat{X}_{i})$, is computed, and compared to a decision threshold $t\in\mathbb{R}^{+}$. The complete workflow of our proposed method to enhance classification is detailed in Algorithm 1, which seeks to approximate the ESC (Definition 3.2). In Appendix B, we discuss the above from an uncertainty perspective, with the aim of providing confidence in the predictions.
@@ 4.2 Visualizing UA-P2E’s ECG Output(s)
In Section 5.3, we provide empirical results regarding the quality of the three visualization approaches mentioned above. These results suggest that the most likely score ECG, as defined in Equation (2), provides the highest quality.
@@ 5 Empirical Study
Unlike prior work, this paper aims to address the uncertainty in the conversion of PPG to ECG signals, and this is obtained by leveraging a diffusion-based conversion methodology. More specifically, we adopt the DDIM framework (Song et al., 2020), aiming to reduce the number of iterations (in our experiments we applied $T=100$ iterations) required for the sampling procedure, and thus increasing the efficiency of our conversion algorithm. The diffusion denoising model is trained using the MIMIC-III matched waveform database (Johnson et al., 2016), which is one of the most extensive dataset for paired PPG and ECG signals. For details on supplementary analysis of overfitting during the training process, we refer the reader to Appendix C.
@@
As for the classification objective, we train and evaluate the classification model, denoted as $f_{\mathcal{X}}$, using the ‘Computing in Cardiology” (CinC) (Reyna et al., 2021) dataset, containing multi-labeled ECG signals of 11 cardiovascular conditions  [^2] The following cardiovascular conditions are considered: (1) Ventricular premature beats, (2) Atrial flutter, (3) Sinus rhythm, (4) Atrial fibrillation, (5) Supraventricular premature beats, (6) Bradycardia, (7) Pacing rhythm, (8) Sinus tachycardia, (9) 1st degree AV block, (10) Sinus bradycardia, and (11) Sinus arrhythmia. that are detectable by lead II ECG signals. Interestingly, our experiments rely in part on a reversed diffusion-based conversion model, designed to sample from the likelihood distribution $\hat{\pi}(Y|X)$, i.e. converting ECG (lead II) to PPG. This reversed sampling is utilized in order to complete the CinC (Reyna et al., 2021) dataset used in our tests, enabling an analysis of challenging cardiac classification tasks. Figure 2 (bottom row) presents typical sampling outcomes of the reversed conversion model. For further details on experimental details, including the datasets used for training and evaluating the conversion and classification models, along with details on data pre-processing and training schemes, we refer the reader to Appendix D.